DontGo:

(Machine C in comments) The table shows the number of candies packed by Machine C. The equation shows the number of candies packed by Machine D. In both representations, x is a measure of the number of minutes and y is a measure of the number of candies packed. Machine D: y = 130x How many more candies could machine D pack than machine C in 11 minutes? 65 130 450 605

DontGo:

DontGo:

Machine C

dude:

Uh you first want to find the equation for the table, do you know how to start?

DontGo:

no..... :(

dude:

Okay, that fine Select any 2 data points in the table

DontGo:

so any 2 ys?? so like 150 and 300??

dude:

A pair 2 x's and their matching y's Sure lets use points (2,150) and (4,300) Okay so subtract the y values

DontGo:

150

DontGo:

right?

dude:

Good and now subtract the x values

DontGo:

2

dude:

Okay (We were using the equation \(\large \frac{y_2-y_1}{x_2-x_1}\) You just subtracted the values so we can now set it up as \(\frac{150}{2}\), can you simplify this?

DontGo:

75

dude:

Good The equation of a line is \(y=mx+b\) We just found the slope \(y=75x+b\)

dude:

Now we can substitute a point back into the equation to find B (Choose a point)

DontGo:

what point?? like give me an example Plzzz

dude:

(4,300)

DontGo:

ok can we go with that?

dude:

Sure \(300=75(4)+b\) Can you solve for B?

DontGo:

if you tell me what I am supposed to do idk sorry

dude:

Isolate B

dude:

What is \(75\times 4\)?

DontGo:

300

dude:

Good and then subtract 300 on both sides \(300=300+b\) \(\color{red}{-300}~~\color{red}{-300}\) ----

DontGo:

wait 300 - 300??

DontGo:

that would be 0

dude:

So b is 0 The second equation is \(y=75x\)

DontGo:

so 130 would be the answer right? y = 130x is the second equation

dude:

Machine C: y=75x Machine D: y=130x To find the difference, first substitute 11 into both equations

DontGo:

130 - 75 = 55 which isn't an answer

dude:

I said substitute 11 into both

DontGo:

what do you mean?

DontGo:

I am soo sorry I am terrible at math

dude:

Change x to 11 and multiply the number with the number in front of it

DontGo:

130 = 1430 75 = 825

dude:

Not subtract those values

dude:

Now*

DontGo:

Thanks

DontGo:

I am going to close this

dude:

Cool np

DontGo:

i clicked best response for u and i am ur fan now too